OpenStudy (anonymous):

Express in the form r sin(x-a) sin x - sqrt(3) cos x (worked solution please, I need to know the method :) )

OpenStudy (raden):

if given an equation : asinx + bsinx, it can be converting to r cos(x- θ) with r = sqrt(a^2 + b^2) to get θ, use defined of tan θ = b/a

OpenStudy (raden):

oppss. i mean if given asinx + bcosx

OpenStudy (raden):

a is the coeff of sinx b is the coeff of cosx

OpenStudy (raden):

now, given sin x - sqrt(3) cos x here a = 1 and b = - 1 so, the value of r = sqrt( 1^2 + (-1)^2) = sqrt(1+1) = sqrt(2)

OpenStudy (raden):

then use defined tanθ = b/a tanθ = -1/1 θ = arc(tan(-1))

OpenStudy (anonymous):

thanks anyway but i worked it out myself :) Thats why i closed it

OpenStudy (raden):

because a has the value +ve(or rechall that the value of sin be (+), it means θ can be in the first quadrant or the 2nd quadrant), while b = -ve (or rechall that the value of cosine be (-), it means θ can be in the second quadrant or the 3rd quadrant) now, most the quadrants for the θ is in the second quadrant. so, θ is 135 degrees. remember tan135 = -1

OpenStudy (raden):

therefore, sin x - sqrt(3) cos x can rewriten becomes sqrt(2) cos(x - 135)

OpenStudy (anonymous):

you are incorrect :(

OpenStudy (raden):

lol which part, i have mistaken

OpenStudy (anonymous):

dont worry man, i got it correct on my own. You dont have to waste your time :P

OpenStudy (raden):

ahhh, you have edited ur equation :P do you want it becomes r cos(x-a) OR r sin(x-a) ?????

OpenStudy (anonymous):

it was always sin(x-a) ?

OpenStudy (raden):

it can be r cos(x-a) OR r sin(x-a) , r tan(x-a), r sec(x-a), r csc(x-a), r cotan(x-a), which one do u like ??? :P

OpenStudy (anonymous):

read the question, it asks for sin(x-a)!

OpenStudy (raden):

before i read it was be r cos(x-a) --"

OpenStudy (anonymous):

yours eyes must have fooled you